OpenStudy (anonymous):

what is 8.5 x 10-3 in standard notation?

OpenStudy (anonymous):

-3 is an exponent

OpenStudy (nurali):

0^(-3) means divide by 10^3 = 1000 so you shift the decimal to the left 5 places 0.0085

OpenStudy (nurali):

Welcome to Openstudy.

OpenStudy (anonymous):

can you help me with another @Nurali

OpenStudy (nurali):

i will try.

OpenStudy (anonymous):

find the balance in the account to the nearest whole dollar 2,500 principal earning 3% compounded annually after 4yrs

OpenStudy (anonymous):

i only have answers of 2,575 2,731, 2,814, 3,350

OpenStudy (nurali):

Formula: A=P*(1+r/n)^n*t A = ? P = $2500, starting amount r = .03 (3% expressed as a decimal) n = 4 (times a year, since it's compounded quarterly) t = 4 (the number of years A=(2500)*(1+0.03/4)^4*4 A=2817.4 I think Third one is Correct